# milliFlow: Scene Flow Estimation on mmWave Radar Point Cloud for Human Motion Sensing

### 12 Jan 2024 v4

@@ 6.3. Sensitivity Analysis
Table 5. Generalization of our model to new activities.
6.4. Generalization to New Activities
@@
In the above experiments, we evaluate our method on the testing set whose subjects are unseen during training. The results demonstrate the generalization ability of our trained model to new users that perform the same ’in-set’ activities. To further test its generalization to new activities, here we evaluate our trained model on the ’out-of-set’ testing set (c.f. Sec. 5.1.2) in which three performed activities are not included in the training set. We can see from Table 5 that, our trained model can still keep an equally good performance, compared to the results in the hallway shown in Table 4, when encountering unseen activities. This demonstrates the ability of our model to cope with new activities in human sensing. As we investigate further we can observe that, on the head bobbing activity, our network achieves the best results across all metrics, achieving a high relax Acc3D of 0.859. We attribute this to the fact that most skeletons are static (except for the neck) when doing this activity (c.f. Fig. 4 (d)), and even for the only moving skeleton, the movement scale is still small. Therefore, estimating the scene flow of points belonging to them is much easier. Similarly, because all skeletons are moving during squatting (c.f. Fig. 4 (d)), it is more challenging for our model to infer scene flow on point clouds captured in this activity. We believe that our trained model can also generalize to other daily human activities.
7. Downstream Task Evaluation
HAR HP
Method oA (%) Gain (%)   mIoU (%)   oA (%) Gain (%)
raw network 47.32 - 49.09 65.75 -
w. strategy 1 57.88 +10.56 52.72 69.27 +3.63 (+3.52)
w. strategy 2 57.78 +10.46 51.04 68.21 +1.95 (+2.46)
Table 6. Evaluation on the HAR and HP task.
@@ 7. Downstream Task Evaluation
As a low-level signal in understanding motions, scene flow can directly enhance low-quality radar point clouds with full per-point displacement information between two frames. Moreover, the latent spatial-temporal representations can be implicitly learned by guiding the network to estimate scene flow, which can be used to support other tasks. Therefore, we envision that learning scene flow estimation can benefit a wide range of higher-level downstream tasks in mmWave-based human sensing. To demonstrate the benefit, here we consider three representative downstream tasks, including HAR, HP and HBPT for evaluation.
7.1. HAR Evaluation
7.1.1. HAR Network.
HAR plays a significant role in a wide range of applications, such as elderly healthcare monitoring (Taylor et al., 2020), smart home (Bianchi et al., 2019) and behaviour surveillance (Khurana and Kushwaha, 2018). The goal of HAR is to identify the specific type of action of a person given a sequence of sensor data. As our focus is to validate the functionality of scene flow to benefit related tasks, here we only design one simple base network for HAR. Given a sequence of $T$ radar point clouds together with their point-level features as input, we follow the process in local feature abstraction (c.f. Sec. 4.3.1) and global feature aggregation (c.f. Sec. 4.3.2) to extract the local-global representations for each point cloud. Then we use another local encoder to extract higher-level features and aggregate the global feature vector again for each frame. We utilize the LSTM network (Hochreiter and Schmidhuber, 1997) to track the temporal relationship across $T$ global vectors and then send the updated hidden state into a final MLP to regress the classification scores $S^{a}=\{s_{i}^{a}\}_{i=1}^{K}$ for $K$ classes. The HAR network can be supervised with a simple cross-entropy loss that measures the difference between $S^{a}$ and a one-hot ground truth activity label.
7.1.2. Scene Flow Application to HAR
We propose two strategies to harness the scene flow network as a plug-and-play module to the downstream network. We believe both can effectively improve downstream performance.
Point cloud decoration. The first way is to directly take the estimated scene flow as point-level raw features and decorate each radar point with them. Such additional scene flow features can explicitly provide the full motion information of radar point clouds. We adopt a two-stage training manner to implement this strategy, which first end-to-end trains the scene flow network and then freezes it to provide raw scene flow features to the downstream network training.
Latent feature recycling. Another way is to leverage the latent representations (c.f. Sec. 4.3.4) encoded by scene flow networks to enhance the low-quality radar point clouds. By guiding the network to estimate accurate scene flow, the learned representation can provide high-level spatial-temporal information to downstream networks in a specific aspect. We follow a joint learning fashion to implement this strategy that takes the final features from the scene flow network as the input features to the downstream network and trains these two networks jointly with a combination of the scene flow and downstream loss functions.
7.1.3. Performance on HAR
We evaluate our proposed two strategies for scene flow application to HAR task on our ”in-set” testing set. The length of each input sequence $T$ for HAR is set as 20 while the number of activity classes $K$ is 5. The evaluation results are shown in Table 6. As we can see, both two proposed strategies can enhance the performance on HAR. Either point cloud decoration (i.e., strategy 1) or latent feature recycling (i.e., strategy 2) yields a remarkable improvement (¿10%) in accuracy, demonstrating their usefulness in helping distinguish between different human actions.
Figure 7. Confusion matrix of HAR.
We also show the breakdown results (with strategy 1) on HAR as the confusion matrix in Fig. 7. As we can see, by decorating radar points with estimated scene flow, our HAR network achieves a high accuracy value on the bowing activity. However, the accuracy values on other activities are not as satisfactory as on bowing. For example, 42% leg swing samples are wrongly predicted as the arm & leg swing activity. This is due to the high similarity between these two activities and in some cases the arms also move with the leg swing, confusing the HAR network.
7.2. HP Evaluation
7.2.1. HP Network.
HP aims to parse human semantic body segments (head, arms, torso, etc.) from sensor data. With radar point clouds as input, our objective of HP is to identify the body parts that correspond to each point. To evaluate the effectiveness of scene flow on this task, we also designed a simple base network for this downstream task. Similar to HAR, HP network also takes sequential radar point clouds as input and utilizes local encoders as well as global aggregation to obtain the latent representation and a global feature vector for each frame. Differently, the human parsing task needs to estimate per-point classification scores for every point rather than the whole sequence. Therefore, after applying the temporal information propagation as in Sec. 4.3.4, we use the point-level final features to regress the parsing scores $S^{p}=\{s_{i}^{p}\}_{i=1}^{V}$ for each point. Here $V$ denotes the number of human body segments. Our HP network is also supervised using a cross-entropy loss with per-point parsing labels. Specifically, we average the loss values over all classes to balance their impact on training. Please see Sec. 5.2.2 for our labelling and denoising process for the HP tasks.
7.2.2. Performance on HP
Figure 8. Human parsing results visualization.
We use the same two strategies proposed in Sec. 7.1.2 to apply scene flow to the HP task on the ”in-set” testing set. The number of body segments for parsing $V$ is 6, including two arms, two legs, head and torso. Leveraging scene flow in two strategies also contributes to better performance on the HP task. Specifically, we observe that directly decorating points with scene flow (i.e., strategy 1) leads to a larger gain than the implicit latent feature recycling (i.e., strategy 2). This is reasonable as our trained scene flow network can provide explicit and reliable per-point full-motion information to downstream tasks, while the learned latent features are implicit and demand further encoding for functioning. Moreover, unlike the two-stage training strategy that only considers one task per time, the scene flow regression and downstream tasks need to be balanced in our joint training fashion, sacrificing their task-specific performance more or less. However, in strategy 2, we only need to run the training once, which is simple to implement in real-world applications. Some examples of our HP results (with strategy 1) are exhibited in Fig. 8. It can be seen that our method can accurately parse radar point clouds into different body parts, close to the ground truth results.
7.3. HBPT Evaluation
7.3.1. Scene flow Application to HBPT
Given the initial position of a part of the human body, our HBPT task aims to track its movement in subsequent frames. Based on our assumption in Sec. 4.4, the human body part can be represented by a human skeleton that is characterized by its two endpoints. As a result, our goal becomes to estimate the transformation matrix of this rigid skeleton given the per-point scene flow estimation, which can be seen as the reverse process of our scene flow labelling. Given point-level scene flow estimation, we first group $N_{t}$ radar points that can be assigned to the skeleton needed to be tracked. Then we generate natural correspondences $\{p_{i},p_{i}+f_{i}\}_{i=1}^{N_{t}}$ using their scene flow and apply the classical Kabsch algorithm (Kabsch, 1976) to solve the rigid skeleton transformation based on them. The new positions of two endpoints can be obtained by applying the transformation to their last positions. By repeating the above steps, we can track a human body part in as many subsequent frames. Note that our method only demands a pre-trained scene flow network and does not need any extra training effort.
7.3.2. Performance on HBPT
Tracking length - mJE (m) $\downarrow$
Activity 1 2 3 4
Arm swing 0.028 0.076 0.097 0.124
Leg swing 0.016 0.071 0.105 0.13
Arm & leg swing 0.030 0.108 0.146 0.178
Average 0.025 0.085 0.116 0.144
@@ 7.3.2. Performance on HBPT
Table 7. Scene flow-based HBPT evaluation results.
@@
We select the sequences of arm swing, leg swing and arm & leg swing activities in the testing set for HBPT evaluation. Specifically, we take the two arms, two legs and both of them, in these three activities respectively as our HBPT targets that are tracked independently across frames. For implementation, we divide the long sequences into short clips with a length of 5 frames and track the body parts within each clip. After initializing the ground truth positions in the first frame, we aim to track each skeleton for the next four frames. The HBPT results at different tracking lengths on three activities are reported in Fig. 7. With accurate scene flow estimation, we can effectively track multiple body parts together for different activities. For one-frame tracking, our method achieves an average mJE of ¡3cm on three activities, demonstrating the capability of scene flow to enable the HBPT task. However, the errors become larger as the tracking length increases, which indicates the happening of the tracking drift. This is inevitable for our method as more radar points associated with other skeletons will be wrongly induced for transformation calculation when the tracking continues. We also observe that the performance on the arm & leg swing activity is worse than the other two activities. This is reasonable as we need to track twice the skeletons in this activity as others, where more skeletons may interfere with each other.
8. Related Work
mmWave Radar-based Human Sensing. The feasibility and versatility of mmWave radar have been extensively demonstrated in various human sensing applications, including vital sign monitoring (Alizadeh et al., 2019; Wang et al., 2020c; Lv et al., 2021), signature verification (Han et al., 2023; Li et al., 2023), fall detection (Jin et al., 2020; Wang et al., 2020a), human tracking and identification (Zhao et al., 2019, 2021; Cui and Dahnoun, 2021; Chen et al., 2023; Gu et al., 2019), gesture recognition (Liu et al., 2021; Hazra and Santra, 2018; Palipana et al., 2021; Liu et al., 2020), activity recognition (Singh et al., 2019; Wang et al., 2021a; Ahuja et al., 2021) and pose estimation, reconstruction (Xue et al., 2021, 2022; Kong et al., 2022). Compared with these applications, our work is unique in that we aim to estimate point-level scene flow vectors instead of providing a holistic output for the whole point cloud. In this way, we can either explicitly augment each radar target with scene flow vectors or implicitly learn robust latent features, which can further benefit a wide range of downstream sensing tasks, such as HAR, HP and HBPT (c.f. Sec. 7).
Scene Flow Estimation on Point Clouds. Recent scene flow works are mostly towards autonomous driving applications and attempt to estimate scene flow on LiDAR point clouds captured from autonomous vehicles. To this goal, different approaches are proposed, including classical methods (Dewan et al., 2016; Pontes et al., 2020; Li et al., 2021) and deep learning-based ones (Liu et al., 2019; Baur et al., 2021; Wu et al., 2020; Wei et al., 2021). Recently, with the advances in point cloud feature learning, deep learning-based methods become more prevalent. According to their learning paradigm, prior works in this thread can be divided into fully-supervised (Wang et al., 2020b, 2021b; Puy et al., 2020; Wei et al., 2021; Behl et al., 2019), self-supervised (Kittenplon et al., 2021; Li et al., 2022; Baur et al., 2021; Mittal et al., 2020; Wu et al., 2020) and weakly-supervised (Gojcic et al., 2021; Dong et al., 2022) learning approaches. Apart from the aforementioned works, our work aims to estimate scene flow for human sensing using mmWave radar. Moreover, our method does not demand any manual annotation efforts, instead, we automatically generate noisy pseudo scene flow labels from corresponding RGB-D images captured by the co-located camera.
Scene Flow Estimation with mmWave Radar. As far as we know, there are limited works that estimate scene flow using mmWave radar. A pioneering work (Ding et al., 2022) proposes a self-supervised learning pipeline for automotive radar scene flow estimation, where three specific loss functions and a two-stage network are bespoken designed for radar point clouds. To improve scene flow performance and enable more downstream applications, a later work (Ding et al., 2023) exploits cross-modal supervision from co-located sensors (e.g. IMU, LiDAR) on modern autonomous vehicles for radar scene flow learning. However, these methods cannot be transferred to human sensing scenarios due to two reasons. First, the radar used for human-centric applications is different from the automotive radar in many aspects, such as measurement properties and their resolution. More importantly, the human object is a non-rigid body, while in autonomous driving, scene dynamics are usually attributed to rigid body motion (e.g. cars and motorcycles) (Gojcic et al., 2021; Dewan et al., 2016; Dong et al., 2022). Such discrepancy in objective indicates different label and supervision generation schemes.
9. Limitation
@@
Despite the progress achieved, there are limitations to be considered in future work. First, we only experiment on a stationary radar platform where only the human subject is movable. This is akin to the scenarios of human sensing in smart spaces and sensors are receded into the background environment. In future work, we plan to evaluate our system on mobile platforms, such as drones and wheeled robots, where the scene flow can be induced by both ego-sensor and object motion. Second, as a proof-of-concept work, the functioning of our method is limited to single subject with coarse skeleton-level body motion (e.g., arm and leg) at this stage. We plan to extend our approach to multi-person scenarios and focus on more fine-grained motion (e.g. finger movement), so that it can support more downstream applications, such as hand gesture recognition and multi-person tracking.
10. conclusion

### 13 Mar 2024 v5

@@ 4.2 Evaluation Setup
Table 2: Breakdown results of our scene flow network.
Table 3: Generalization of our model to new activities.
4.3 Scene Flow Evaluation
Compared to the State of the Arts. We first compare our scene flow network with state-of-the-art methods of point-based scene flow estimation. Our baselines include 7 deep learning-based methods [52, 83, 43, 60, 81, 17, 23] whose networks are also supervised with pseudo scene flow labels, and one non-learning-based method Graph Prior [59] that solves scene flow via online optimization. The evaluation results on the ‘in-set’ testing set can be seen in Table 3. The scene flow network of milliFlow achieves a cm-level average EPE3D (i.e., 4.6cm) and a high relax Acc3D of 70.3%, ranking 1st on all metrics with a large margin compared to the baselines. This satisfactory performance proves the efficacy of our method to address the challenges existing in our scene flow task. Note that our network is trained with pseudo labels automatically generated using RGB-D images (cf. Sec. 3.5), which does not demand any manual annotation efforts.
Qualitative Results. Our output examples, illustrated in Fig. 5, effectively demonstrate our method’s ability to produce reliable scene flow for diverse activities and subjects across three environments. Despite the inherent sparsity and noise in radar point clouds, our network adeptly learns representative features through global feature aggregation and robust scene flow vector regression. Notably, even when some body parts are absent in the radar data, our approach compensates by utilizing historical information, thereby benefiting the current scene flow estimation and ensuring robust performance across successive frames.
Runtime Efficiency. We test the runtime efficiency of our scene flow network on a single NVIDIA RTX 3090 GPU. Given sequential testing radar point clouds, we feed them one by one into our trained model for inference. As a result, our model has real-time performance with one inference step in 74ms ($\sim$13.5Hz). Moreover, the maximum allocated GPU memory is only 134 MB during inference. This minimal memory usage enables our network to operate in parallel with downstream networks, underscoring its practicality for real-time applications.
Figure 5: Qualitative scene flow results. radar points and scene flow vectors are projected onto the image and the colour red is used for the ground truth while blue for ours.
Ablation Study. To validate the effectiveness of key components within our scene flow network, we systematically disabled each, i.e., temporal propagation (TP), global aggregation (GA), context feature (CF), and constrained regression (CR), and observed their effects on performance, as presented in Table 3. Overall, the full version of our network (row (a)) yields the best results and each component helps to elevate the performance on each metric (from row (e) to (a)). By utilizing information from previous frames, temporal propagation significantly improved scene flow estimation, achieving a 13.2% reduction in EPE3D and a 6.3% increase in strict Acc3D accuracy. Global aggregation, employing an attention mechanism, bolstered local features with global context, leading to a 13.1% improvement in EPE3D. The inclusion of context features into our flow embedding, aimed at retaining source frame context, provided a modest but surprising boost, lowering the overall EPE3D by 0.01m. Lastly, the biggest improvement (i.e., a 0.012m decrease on EPE3D) is brought by our constrained regression, in which the scene flow component on each axis is clamped by a fixed threshold (e.g. 0.1m). This is reasonable as non-viable results, for example, a scene flow vector with a length of 0.5m, can seriously degrade our results.
@@ 4.3 Scene Flow Evaluation
Generalization to New Activities. In the above experiments, we evaluate our method on the testing set whose subjects are unseen during training. The results demonstrate the generalization ability of our trained model to new users that perform the same ‘in-set’ activities. To further test its generalization to new activities, here we evaluate our trained model on the ‘out-of-set’ testing set in which three performed activities are not included in the training set. We can see from Tab. 3 that, our trained model can still keep an equally good performance when encountering unseen activities. This demonstrates the ability of our model to cope with new activities in human sensing. We also observed that the performance to new activities is better than the overall performance shown in Tab. 3. This is because our subjects either keep most of their bodies static (i.e., heading bobbing) or stay completely static in many frames (i.e., sitting and squatting) when doing unseen activities. As a result, estimating the scene flow of points belonging to them is much easier. We believe that our trained model can also generalize to other daily human activities.
4.4 Downstream Task Evaluation
Table 4: Evaluation on the benefit of scene flow for the HAR task.
Table 5: Evaluation on the benefit of scene flow for the HP task.
Method Raw w. S1 Gain w. S2 Gain
Ours 47.32 57.88 +10.56 57.78 +10.46
MMPointGNN [29] 52.46 60.16 +7.70 59.94 +7.48
RadHAR [67] 44.65 49.98 +5.33 50.53 +5.88
Average 48.14 56.01 +7.87 56.08 + 7.94
Method mIoU (%) Gain (%) oA (%) Gain (%)
Raw 49.09 - 65.75 -
w. S1 52.72 +3.63 69.27 +3.52
w. S2 51.04 +1.95 68.21 +2.46
Table 5: Evaluation on the benefit of scene flow for the HP task.
@@ 4.4 Downstream Task Evaluation
As a low-level signal in understanding motions, scene flow can directly enhance low-quality radar point clouds with full per-point displacement information between two frames. Moreover, the latent spatial-temporal representations can be implicitly learned by guiding the network to estimate scene flow, which can be used to support other tasks. Therefore, we envision that learning scene flow estimation can benefit a wide range of higher-level downstream tasks in mmWave-based human sensing. To demonstrate the benefit, here we consider three representative downstream tasks, including human activity recognition (HAR), human parsing (HP) and human body part tracking (HBPT) for evaluation.
HAR Evaluation Setup. HAR plays a significant role in a wide range of applications, such as elderly healthcare monitoring [69], smart home [12] and behaviour surveillance [41]. To validate the functionality of scene flow to benefit HAR, we design a HAR base network by combining some key components from Sec. 3.4 and also utilizing the LSTM network [34] to track the temporal relationship before regressing the classification scores. Besides, two state-of-the-art methods [29, 67] bespoken for mmWave-based HAR are selected for a more persuasive comparison. Particularly, we design two strategies to harness the scene flow network as a plug-and-play module to the HAR network. Strategy 1 (S1) directly takes the estimated scene flow as point-level raw features and decorates each radar point with them, while Strategy 2 (S2) leverages the latent representations encoded by scene flow networks to enhance the low-quality radar data. Please see more implementation details in supplementary materials.
HAR Evaluation Results. We evaluate our proposed two strategies for scene flow application to HAR task on our ‘in-set’ testing set. The length of each input sequence $T$ for HAR is set as 20. The evaluation results are shown in Table 5. As we can see, both two proposed strategies can enhance the performance of our base network and state-of-the-art methods on HAR, demonstrating their usefulness in helping distinguish between different human activities. It can be also observed that the average accuracy shown in Tab. 5 is relatively low compared to those reported in [29, 67]. We credit this to our subject activities being very similar and inter-included to some extent. For example, the arm & leg swing activity can be separated into arm swing and leg swing, which are also inside our ‘in-set’ activities. As a result, both ours and the two states of the arts find it more difficult to correctly distinguish between these activities. The breakdown results (ours with S1) on HAR as the confusion matrix is shown in Fig. 6. As we can see, our HAR network achieves a relatively high accuracy value on the bowing activity since it has apparently different motion patterns from others. However, the accuracy values on other activities are not as satisfactory as on bowing. For example, 42% leg swing samples are wrongly predicted as the arm & leg swing activity. This further justifies our explanation that some activities are so similar and thus can easily confuse our HAR networks in classification.
Figure 6: Confusion matrix of HAR.
@@
Table 6: Scene flow-based HBPT evaluation results. Tracking length - mJE (m) $\downarrow$ Activity 1 2 3 4 Arm swing 0.028 0.076 0.097 0.124 Leg swing 0.016 0.071 0.105 0.130 Arm & leg swing 0.030 0.108 0.146 0.178 Average 0.025 0.085 0.116 0.144
Figure 6: Confusion matrix of HAR.
HP Evaluation Human parsing aims to parse human semantic body segments (head, arms, torso, etc.) from sensor data. With radar point clouds as input, our objective of HP is to identify the body parts that correspond to each point. To evaluate the effectiveness of scene flow on this task, we also designed a base network for this downstream task using key components from Sec. 3.4. Both two scene flow application strategies proposed for the HAR above are also used here for the HP task. The number of body segments for parsing is 6, encompassing two arms, two legs, head and torso. As seen in Tab. 5, leveraging scene flow in two strategies also contributes to better performance on the HP task. Directly applying scene flow to points (S1) yields better results than using latent feature recycling (S2), as the former provides explicit per-point motion information. In contrast, latent features are less direct and require additional processing. Some examples of our HP results (with Strategy 1) are exhibited in Fig. 7. Thanks to the enhancement by the scene flow, our method can accurately parse radar point clouds into different body parts, close to the ground truth results.
Figure 7: Human parsing results visualization.
HBPT Evaluation. Given the initial position of a human body part, our human body part tracking task aims to track its movement in subsequent frames. With point-level scene flow estimation, we first group $N_{t}$ radar points that can be assigned to the skeleton needed to be tracked (cf. Fig. 3). Then we generate natural correspondences $\{p_{i},p_{i}+f_{i}\}_{i=1}^{N_{t}}$ using their scene flow and apply the classical Kabsch algorithm [40] to solve the rigid skeleton transformation based on them. This updates the endpoints’ positions, allowing for tracking of the human body part across frames without additional training, relying only on a pre-trained scene flow network. For evaluation, we select the sequences of arm swing, leg swing and arm & leg swing activities in the testing set and take the two arms, legs and both of them respectively as our tracking targets.
@@
For implementation, we divide the long sequences into short clips with a length of 5 frames and track the body parts within each clip. After initializing the ground truth positions in the first frame, we aim to track each skeleton for the next four frames. The HBPT results at different tracking lengths on three activities are reported in Fig. 6. With accurate scene flow estimation, we can effectively track multiple body parts together for different activities. For one-frame tracking, our method achieves an average mJE of <3cm on three activities, demonstrating the capability of scene flow to enable the HBPT task. However, the errors become larger as the tracking length increases, which indicates the happening of the tracking drift. This is inevitable for our method as more radar points associated with other skeletons will be wrongly induced for transformation calculation when the tracking continues. We also observe that the performance on the arm & leg swing activity is worse than the other two activities. This is reasonable as we need to track twice the skeletons in this activity as others, where more skeletons may interfere with each other during tracking.
5 Limitation and Future Work
@@
Despite the progress achieved, there are limitations to be considered in future work. First, we only experiment on a stationary radar platform where only the human subject is movable. This is akin to the scenarios of human motion sensing in smart spaces and sensors are receded into the background environment. In future work, we plan to evaluate our system on mobile platforms, such as drones and wheeled robots, where the scene flow can be induced by both ego-sensor and subject motion. Second, as a proof-of-concept work, the functioning of our method is restricted to the single subject with face-forward body motion at this stage. We plan to extend our approach to multi-person scenarios and introduce variations in the subjects’ position and orientation relative to the sensor in the next study. Lastly, the number of activities included in our dataset is limited as we only design activities that can be separated into the rigid motion of multiple body skeletons. This fits our need to annotate scene flow labels based on body skeleton displacement. In our upcoming work, we will design more complex and smaller activities and exploit a novel approach to derive the scene flow.
6 Conclusion